Exercise 4 Alpha Corp oration reports the following information at the end of 2018: Accumulated depreciation Cash and bank accounts Common stock Discount on bond payable Dividends distributed in 2018 Income tax payable Interest expenses Net income 2018 Number of common shares issued Number of preferred shares issued Paid in capital in excess of par-common Paid in capital in excess of par - preferred Preferred dividends in arrears 415,000 28,550 850,000 64,800 18,000 6,800 5,600 35,0001- 1,700,000 300,000 900,000 200,000 6,000 150,000 6,800 1,223,000 1,524,000 120,000 Preferred stock, 5% Prepaid rent Retained earnings (beginning balance, credit) Sales revenue Treasury shares (common stock, unit price 10) Unearned revenue 18,500 Choosing only the appropriate data among those listed above, determine the fol clearly showing the calculations performed: a) Total contributed (paid-in) capital b) Retained earnings c) Shareholders equity d) Par value per share of common stock e) Par value per share of preferred stock f) Number of common and preferred shares outstanding g) Book value per share of common stock h) Book value per share of preferred stock
Answer #1

Part A

Total contributed (paid-in) capital = common stock + paid in capital in excess of par – common + preferred stock + paid in capital in excess of par – preferred = 850000+900000+150000+200000 = 2100000

Part B

Retained earnings = retained earnings, beginning balance + net income – dividends = 1223000+35000-18000 = 1240000

Part C

Shareholder’s equity = total paid in capital + retained earnings – treasury shares = 2100000+1240000-120000 = 3220000

Part D

Par value per share of common stock = common stock / number of common shares issued = 850000/1700000 = $0.5 per share

Part E

Par value per share of preferred stock = preferred stock / number of preferred shares issued = 150000/300000 = $0.5 per share

Part F

Number of common share outstanding = 1700000-(120000/10) = 1688000 shares

Number of preferred share outstanding = 300000 shares

Part G

Book value per share of common stock = (common stock + paid in capital in excess of par – common-treasury shares)/ Number of common share outstanding = (850000+900000-120000)/1688000 = $0.97

Part H

Book value per share of preferred stock = (preferred stock + paid in capital in excess of par – preferred)/ Number of preferred share outstanding = (150000+200000)/300000 = $1.17 per share